Question 34 Nicole Limited is a company that produces machinery to customer orders, using a normal job-order cost system. It applies manufacturing overhead to production using a predetermined rate. This overhead rate is set at the beginning of each fiscal year by forecasting the years overhead and relating it to direct labour costs. The budget for 2020 was as follows: Direct labour $1,800,000 900,000 Manufacturing overhead As at the end of the year, two jobs were incomplete. These were 1768B, with total direct labour charges of $110,000, and 1819C, with total direct labour charges of $390,000. On these jobs, machine hours were 287 hours for 17688 and 647 hours for 1819C. Direct materials issued for 1768B amounted to $220,000, and for 1819C they amounted to $420,000 Total charges to the Manufacturing Overhead Control account for the year were $897,000, and direct labour charges made to all jobs amounted to $1,573,600, representing 247,200 direct labour hours. There were no beginning inventories. In addition to the ending work in process just described, the ending finished goods inventory account showed a balance of $556,000 Sales for the year amounted to $6,201,300; cost of goods sold totalled $3,614,000; and sales, general, and administrative expenses were $1,847,800 The above amounts for inventories and the cost of goods sold have not been adjusted for any over- or under-application of manufacturing overhead to production. It is the companys practice to allocate any over- or under-applied overhead to inventories and the cost of goods sold

Your answer is partially correct. Try again. Prorate the amount calculated in part (a) based on the ending balances (before prorating) of Work in Process, Finished Goods, and Cost of Goods Sold. (Round allocation percentage to 2 decimal places, e.g. 15.25% and final answers to the nearest whole dollar, eg, 5,275.) Allocation Cost of goods sold 75,002 Finished goods inventory 11,539 WIP inventory 23,659 Total 110,200
Answer #1
a)Overhead rate =Overhead /Direct labor cost estimated
                  = 900000/1800000
                  = .50 or50 % of direct labor cost
Actual overhead : 897,000
Applied overhead : Actual direct labor cost *overhead rate
                = 1573600*.50 786800
overhead underapplied =Actual -applied
              = 897000 - 786800 110200
b) Work in process : cost of job1768B + Cost of job1819c
         = [220000+110000+(110000*.50)]+[420000+390000+(390000*.50)]
         = 385000+1005000 1390000
i ii iii=110200*ii
Amount % of allocation underapplied Overhead allocation
Work in process 1390000 25.00% 27550
Finished goods inventory 556000 10.00% 11020
Cost of goods sold 3614000 65.00% 71630